# On Malle’s conjecture for the product of symmetric and nilpotent groups

### v2

@@ 1.1. Motivation and historical context
Let $k$ be a number field and let $n\in\mathbb{Z}_{\geq 1}$. Let $\mathcal{G}$ be a transitive subgroup of $S_{n}$. Given a number field extension $K/k$, denote by $\widetilde{K}$ its Galois closure over $k$. Suppose that $[K\mathrel{\mathop{:}}k]=n$. We enumerate the embeddings $\iota_{j}\mathrel{\mathop{:}}K\hookrightarrow\bar{K}$ for $j=1,\dots,n$, and for $\sigma\in\mathcal{G}$, consider the composite
@@
$\operatorname{Disc}(K)\mathrel{\mathop{:}}=|\operatorname{Norm}_{k/\mathbb{Q}}(\operatorname{disc}(K/k))|.$
For $X>0$, let $N_{k}(\mathcal{G};X)$ be the number of extensions $K/k$ with $[K\mathrel{\mathop{:}}k]=n$ for which $\operatorname{Gal}(\widetilde{K}/k)$ is isomorphic to $\mathcal{G}$ as a permutation subgroup of $S_{n}$, and $\operatorname{Disc}(K)\leq X$. The quantity $N_{k}(\mathcal{G};X)$ depends not only on $\mathcal{G}$ and $k$, but also on the embedding of $\mathcal{G}$ in $S_{n}$, and this is suppressed in our notation. For $g\in S_{n}$, set
(1.1) $\operatorname{ind}(g)\mathrel{\mathop{:}}=n-\text{number of orbits for }g.$
We set $1=1_{\mathcal{G}}$ to denote the identity element in $\mathcal{G}$ and $\mathcal{G}^{\ast}\mathrel{\mathop{:}}=\mathcal{G}\backslash\{1\}$.
@@
$a(\mathcal{G})\mathrel{\mathop{:}}=\left(\operatorname{min}\{\operatorname{ind}(g)\mid g\in\mathcal{G}^{\ast}\}\right)^{-1}.$
The absolute Galois group $G_{k}\mathrel{\mathop{:}}=\operatorname{Gal}(\overline{\mathbb{Q}}/k)$ of $k$ acts on the set of conjugacy classes $C(\mathcal{G})$ of $\mathcal{G}$ via the action of $\overline{\mathbb{Q}}$-characters of $\mathcal{G}$. We define
$b(k,\mathcal{G})\mathrel{\mathop{:}}=\#\{C\in C(\mathcal{G})\mathrel{\mathop{:}}\operatorname{ind}(C)=\operatorname{ind}(\mathcal{G})\}/G_{k}.$
The following conjecture (cf. [Mal04, Conjecture 1.1]) is referred to as the strong form of Malle’s conjecture
@@
  • For finite nilpotent groups $G$ sitting in $S_{|G|}$, satisfying additional conditions, the conjecture was recently resolved by Koymans and Pagano [KP23]. The weak version of the conjecture for all finite nilpotent groups was previously settled by Klüners and Malle [KM04].
@@
    The conditions on $A$ were subsequently relaxed due to work of Masri, Thorne, Wei-Lun and Wang [MTTW20].
@@ 1.2. Main result
Let $G$ be a finite nilpotent group and $\ell_{G}$ the smallest prime factor of $|G|$. We recall that $a(G)=\frac{\ell_{G}}{(\ell_{G}-1)|G|}$, $a(S_{n})=1$, and $b(k,S_{n})=1$ for all $n\geq 2$. It is easy to see that $a(S_{n}\times G)=\frac{1}{|G|}$. Since ${a(S_{n})}{n}>{a(G)|G|}$ we deduce that $b(k,S_{n}\times G)=b(k,S_{n})$. We now state our main result.
@@
Let $G$ be a non-trivial finite nilpotent group and $k$ a number field. Consider the regular representation $\operatorname{reg}_{G}\mathrel{\mathop{:}}G\hookrightarrow S_{|G|}$. Here, $S_{n}\times G$ embeds naturally in $S_{n|G|}$ via natural inclusions
@@
Then
@@ 1.3. Outline of the proof
We closely follow the method in [Wan21]. Two complications have to be addressed. First, the method requires strong uniformity estimates for nilpotent groups. We address this in Section 3. It is worth noting that although a weaker uniformity estimate suffices for proving the main theorem, we pursue a sharp bound to attain the best possible uniformity estimate. This result stands as a potentially independent point of interest. Second, we do not have precise asymptotics for counting $G$-extensions satisfying prescribed local conditions. For more details, see Section 4.
@@ Acknowledgments
The authors would like to thank Peter Koymans and Jiuya Wang for their insightful comments and suggestions.
@@ 2. Preliminary results
We begin this section by recalling some notation. Throughout, $k$ will be a fixed number field. Set $\Omega_{k}$ to denote the set of all primes of $k$. For any non-zero ideal $\mathfrak{I}\subset\mathcal{O}_{k}$, set $|\mathfrak{I}|\mathrel{\mathop{:}}=|\operatorname{Norm}_{k/\mathbb{Q}}(\mathfrak{I})|$. Given any finite extensions $K/k$ we denote by $\mathrm{disc}(K/k)$ the relative discriminant ideal in $k$ and set $\mathrm{Disc}(K)\mathrel{\mathop{:}}=|\mathrm{disc}(K/k)|$. We use $\mathfrak{p}$ to denote a finite prime in $k$. Take $\mathrm{disc}_{\mathfrak{p}}(K/k)$ to be the ideal $\mathfrak{p}^{\operatorname{ord}_{\mathfrak{p}}(\mathrm{disc}(K/k))}$ and $\mathrm{Disc}_{\mathfrak{p}}(K)$ its absolute norm. We denote the inertia group at a prime $\mathfrak{p}$ by $I_{\mathfrak{p}}(K)$. Let $G$ be a finite nilpotent group and $n\in\{3,4,5\}$ be an integer. Note that $G$ sits inside $S_{|G|}$ via the regular representation. Therefore, we view the product $S_{n}\times G$ as a subgroup of $S_{n}\times S_{|G|}\subset S_{n|G|}$.
@@
Let $K/k$ and $L/k$ be number field extensions with $n\mathrel{\mathop{:}}=[K\mathrel{\mathop{:}}k]$ and $m\mathrel{\mathop{:}}=[L\mathrel{\mathop{:}}k]$. Assume that $[KL\mathrel{\mathop{:}}k]=[K\mathrel{\mathop{:}}k][L\mathrel{\mathop{:}}k]$. Then we have that
@@
If $K/k$ and $L/k$ are tamely ramified at $\mathfrak{p}$, then the inertia groups at $\mathfrak{p}$ are cyclic. Let $n\mathrel{\mathop{:}}=[K\mathrel{\mathop{:}}k]$ and $m\mathrel{\mathop{:}}=[L\mathrel{\mathop{:}}k]$. Let $g_{K}\in S_{n}$ and $g_{L}\in S_{m}$ respectively denote generators for these groups. Then, $\mathrm{disc}_{\mathfrak{p}}(K)=p^{\mathrm{ind}(g_{K})}$ and $\mathrm{disc}_{\mathfrak{p}}(L)=p^{\mathrm{ind}(g_{L})}$, where $\mathrm{ind}(g)$ denotes the index of $g$ (cf. (1.1)). The next result is used to determine $\mathrm{disc}_{\mathfrak{p}}(KL)$ in terms of $\mathrm{disc}_{\mathfrak{p}}(K)$ and $\mathrm{disc}_{\mathfrak{p}}(L)$ at the tamely ramified places. Denote by $\widetilde{K}$ (resp. $\widetilde{L}$) the Galois closure of $K$ (resp. $L$).
@@
    Here $n=[K\mathrel{\mathop{:}}k]$ and $m=[L\mathrel{\mathop{:}}k]$.
@@
Since $(|c_{i}|,|d_{j}|)=1$ for all $i,j$ the element $(c_{i},d_{j})\in S_{nm}$ is a cycle of length $|c_{i}||d_{j}|$ for each $i,j$. These cycles in $S_{nm}$ are disjoint from each other and $(\sigma,\tau)=\prod_{(i,j)}(c_{i},d_{j})$. Hence, the number of cycles in the decomposition of $(\sigma,\tau)$ is the product of the number of cycles in the decomposition of $\sigma$ and $\tau$. We recall that the number of such cycles is $n-\mathrm{ind}(\sigma)$ for $\sigma\in S_{n}$ and $m-\mathrm{ind}(\tau)$ for $\tau\in S_{m}$. Therefore,
@@
Let $\ell_{1},\dots,\ell_{c}$ be the primes that divide $|G|$ and denote by $G(\ell_{i})$ the $\ell_{i}$-Sylow subgroup of $G$. Since $G$ is a nilpotent group, we have the product decomposition $G\mathrel{\mathop{:}}=\prod_{i=1}^{c}G(\ell_{i})$. Denote by $\ell_{G}$ the smallest prime factor of $|G|$. Given $\sigma\in S_{n}$ and $g\in G$ we compute the index of $(\sigma,g)\in S_{n|G|}$ using the Lemma 2.4. The method employed by Wang for abelian groups, as detailed in [Wan21, Lemmas 2.5, 2.6 and 2.7] generalizes to nilpotent groups. For the sake of completeness, we provide complete details. In what follows, we set $G^{\ast}$ to denote $G\backslash\{1_{G}\}$ and $\operatorname{ind}(G)\mathrel{\mathop{:}}=\operatorname{min}\{\operatorname{ind}(g)\mid g\in G^{\ast}\}$.
@@
Let $G$ be a finite nilpotent group.
@@
    $\mathrm{ind}((12),g)/|G|>2\text{ and }\mathrm{ind}((123),g)/|G|>1$
@@
    $\mathrm{ind}((12),g)/|G|>2\text{ and }\mathrm{ind}((12)(34),g)/|G|>1,$
    $\mathrm{ind}((123),g)/|G|>3\text{ and }\mathrm{ind}((1234),g)/|G|>2$
@@
    $\mathrm{ind}(\sigma,g)/|G|\geq 1+\mathrm{ind}(\sigma)-1/7$
@@
    $\mathrm{ind}(\sigma,g)/|G|\geq 12/7+\mathrm{ind}(\sigma).$
@@
We begin with (1). For $(12)\in S_{3}$, using Lemma 2.4 we deduce that for all $g\in G$ we have that $\mathrm{ind}((12),g)=|G|+2\cdot\mathrm{ind}(g)$ as $\mathrm{ind}(12)=1$. We deduce that $\mathrm{ind}((12),g)/|G|\geq 1+2\cdot\mathrm{ind}(G)/|G|>2$ since $\ell_{G}>2$.
For $(123)\in S_{3}$, if $3\nmid|G|$, then again using Lemma 2.4 we obtain that for all $g\in G$, $\mathrm{ind}((123),g)=2|G|+\mathrm{ind}(g)$ as $\mathrm{ind}(123)=2$. Hence the inequality holds in this case. If $3\mid|G|$, then $G=G(3)\times\widetilde{G}$. Given any $g\in G$ we can write it as $(g(3),\widetilde{g})\in S_{|G(3)|}\times S_{\widetilde{G}}\subset S_{|G|}$. Further we note that $\mathrm{ind}((123),g)=\mathrm{ind}((123),(g(3),\widetilde{g}))=\mathrm{ind}(((123),g(3)),\widetilde{g})$ in $S_{|G|}$. Since $3\nmid|\widetilde{G}|$ we can use Lemma 2.4 to deduce that $\mathrm{ind}(((123),g(3)),\widetilde{g})=\mathrm{ind}((123,g(3))\cdot(|\widetilde{G}|-\mathrm{ind}(\widetilde{g}))+3\cdot|G(3)|\cdot\mathrm{ind}(\widetilde{g})$. Now, since the smallest possible value of $\mathrm{ind}((123,g(3))$ is $2|G(3)|$ when $g(3)$ is identity and smallest possible value for $\mathrm{ind}(\widetilde{g})$ is $\mathrm{ind}(\widetilde{G})$ by definition. Hence, $\mathrm{ind}((123,g(3))\cdot(|\widetilde{G}|-\mathrm{ind}(\widetilde{g}))+3\cdot|G(3)|\cdot\mathrm{ind}(\widetilde{g})\geq 2|G|+|G(3)|\cdot\mathrm{ind}(\widetilde{G})>|G|$. This concludes the proof.
@@
(3) Using Lemma 2.4 we conclude that $\mathrm{ind}(\sigma,g)=\mathrm{ind}(\sigma)|G|+5\cdot\mathrm{ind}(g)-\mathrm{ind}(\sigma)\cdot\mathrm{ind}(g)\geq\mathrm{ind}(\sigma)\cdot|G|+(5-\mathrm{ind}(\sigma))\cdot\mathrm{ind}(g)\geq\mathrm{ind}(\sigma)\cdot|G|+(5-\mathrm{ind}(\sigma))\cdot\mathrm{ind}(G)\geq\mathrm{ind}(\sigma)\cdot|G|+(5-\mathrm{ind}(\sigma))\cdot(\ell_{G}-1)|G|/\ell_{G}\geq\mathrm{ind}(\sigma)\cdot|G|+(5-\mathrm{ind}(\sigma))\cdot 6|G|/7$ since $\ell_{G}\geq 7$. For $\sigma\in S_{5}$ not conjugate to $(12345)$ we have that $(5-\mathrm{ind}(\sigma))\geq 2$ and hence $\mathrm{ind}(\sigma)\cdot|G|+(5-\mathrm{ind}(\sigma))\cdot 6|G|/7\geq\mathrm{ind}(\sigma)\cdot|G|+12|G|/7$. This concludes the proof of Proposition. ∎

### v3

@@ 1.1. Motivation and historical context
[…]
@@
$\operatorname{Disc}(K):=|\operatorname{Norm}_{k/\mathbb{Q}}(\operatorname{disc}(K/k))|.$
[…]
(1.1) $\operatorname{ind}(g):=n-\text{number of orbits for }g.$
We set $1=1_{\mathcal{G}}$ to denote the identity element in $\mathcal{G}$ and $\mathcal{G}^{\ast}:=\mathcal{G}\backslash\{1\}$.
@@
$a(\mathcal{G}):=\left(\operatorname{min}\{\operatorname{ind}(g)\mid g\in\mathcal{G}^{\ast}\}\right)^{-1}.$
The absolute Galois group $G_{k}:=\operatorname{Gal}(\overline{\mathbb{Q}}/k)$ of $k$ acts on the set of conjugacy classes $C(\mathcal{G})$ of $\mathcal{G}$ via the action of $\overline{\mathbb{Q}}$-characters of $\mathcal{G}$. We define
$b(k,\mathcal{G}):=\#\{C\in C(\mathcal{G}):\operatorname{ind}(C)=\operatorname{ind}(\mathcal{G})\}/G_{k}.$
The following conjecture (cf. [Mal04, Conjecture 1.1]) is referred to as the strong form of Malle’s conjecture.
@@
  • Let $G$ be a finite nilpotent group, $p$ be the smallest prime divisor of $|G|$. Assume that all elements of order $p$ are central. Consider $G$ sitting in $S_{|G|}$ via the regular representation. The strong form of Malle’s conjecture in this context was recently resolved by Koymans and Pagano [KP23]. The weak version of the conjecture for all finite nilpotent groups $G$ sitting in $S_{|G|}$ via the regular representation, was previously settled by Klüners and Malle [KM04].
@@
    The conditions on $A$ were subsequently removed due to work of Masri, Thorne, Wei-Lun and Wang [MTTW20].
@@ 1.2. Main result
Let $G$ be a finite nilpotent group and $\ell_{G}$ the smallest prime factor of $|G|$. We view $G$ as a subgroup of $S_{|G|}$ via the regular representation. On the other hand, $S_{n}$ is in its natural degree $n$ representation (thus contained in $S_{n}$ via the identity map). We recall that $a(G)=\frac{\ell_{G}}{(\ell_{G}-1)|G|}$, $a(S_{n})=1$, and $b(k,S_{n})=1$ for all $n\geq 2$. It is easy to see that $a(S_{n}\times G)=\frac{1}{|G|}$. Since ${a(S_{n})}{n}>{a(G)|G|}$ we deduce that $b(k,S_{n}\times G)=b(k,S_{n})$, cf. [Wan21, p.91 l.-3 – p. 92, l.2]. We now state our main result.
@@
Let $G$ be a non-trivial finite nilpotent group and $k$ a number field. Consider the regular representation $\operatorname{reg}_{G}:G\hookrightarrow S_{|G|}$. Here, $S_{n}\times G$ embeds in $S_{n|G|}$ via natural inclusions
@@
Then there exists a constant $c(k,S_{n}\times G)>0$ such that
@@ 1.3. Outline of the proof
We closely follow the method in [Wan21]. Two complications have to be addressed. First, the method requires strong uniformity estimates for nilpotent groups. We address this in Section 3. It is worth noting that although a weaker uniformity estimate suffices for proving the main theorem, we pursue a sharp bound to attain the best possible uniformity estimate (see Proposition 3.6). This result stands as a potentially independent point of interest. Proposition 4.6 gives us a crucial ingredient for proving precise asymptotics for $S_{n}\times G$-extensions. This technical result is the analogue of [Wan21, p. 117 (5.6)], however, it is proven via a different strategy. For more details, see Section 4.
@@ Acknowledgments
We would like to thank Peter Koymans and Jiuya Wang for their insightful comments and suggestions. We also extend our sincerest thanks to the anonymous referee for the excellent and timely report.
@@ 2. Preliminary results
We begin this section by recalling some notation. Throughout, $k$ will be a fixed number field. Set $\Omega_{k}$ to denote the set of all primes of $k$. For any non-zero ideal $\mathfrak{I}\subset\mathcal{O}_{k}$, set $|\mathfrak{I}|:=|\operatorname{Norm}_{k/\mathbb{Q}}(\mathfrak{I})|$. Given any finite extensions $K/k$ we denote by $\mathrm{disc}(K/k)$ the relative discriminant ideal in $k$ and set $\mathrm{Disc}(K):=|\mathrm{disc}(K/k)|$. We use $\mathfrak{p}$ to denote a finite prime in $k$. Take $\mathrm{disc}_{\mathfrak{p}}(K/k)$ to be the ideal $\mathfrak{p}^{\operatorname{ord}_{\mathfrak{p}}(\mathrm{disc}(K/k))}$ and $\mathrm{Disc}_{\mathfrak{p}}(K)$ its absolute norm. We denote the inertia group at a prime $\mathfrak{p}$ by $I_{\mathfrak{p}}(K)$. Let $G$ be a finite nilpotent group and $n\in\{3,4,5\}$ be an integer. Note that $G$ sits inside $S_{|G|}$ via the regular representation. Therefore, we view the product $S_{n}\times G$ as a subgroup of $S_{n}\times S_{|G|}\subset S_{n|G|}$.
@@
[…]
@@
If $K/k$ and $L/k$ are tamely ramified at $\mathfrak{p}$, then the inertia groups at $\mathfrak{p}$ are cyclic. Let $n:=[K:k]$ and $m:=[L:k]$. Let $g_{K}\in S_{n}$ and $g_{L}\in S_{m}$ respectively denote generators for these groups. Then, $\mathrm{disc}_{\mathfrak{p}}(K)=p^{\mathrm{ind}(g_{K})}$ and $\mathrm{disc}_{\mathfrak{p}}(L)=p^{\mathrm{ind}(g_{L})}$, where $\mathrm{ind}(g)$ denotes the index of $g$ (cf. (1.1)). The next result is used to determine $\mathrm{disc}_{\mathfrak{p}}(KL)$ in terms of $\mathrm{disc}_{\mathfrak{p}}(K)$ and $\mathrm{disc}_{\mathfrak{p}}(L)$ at the tamely ramified places. Denote by $\widetilde{K}$ (resp. $\widetilde{L}$) the Galois closure of $K$ (resp. $L$).
@@
    Here $n=[K:k]$ and $m=[L:k]$.
@@
The result is essentially the same as [Wan21, Theorem 2.2]. Since the proof is relatively short, we include it for completeness. Since $(|c_{i}|,|d_{j}|)=1$ for all $i,j$ the element $(c_{i},d_{j})\in S_{nm}$ is a cycle of length $|c_{i}||d_{j}|$ for each $i,j$. These cycles in $S_{nm}$ are disjoint from each other and $(\sigma,\tau)=\prod_{(i,j)}(c_{i},d_{j})$. Hence, the number of cycles in the decomposition of $(\sigma,\tau)$ is the product of the number of cycles in the decomposition of $\sigma$ and $\tau$. We recall that the number of such cycles is $n-\mathrm{ind}(\sigma)$ for $\sigma\in S_{n}$ and $m-\mathrm{ind}(\tau)$ for $\tau\in S_{m}$. Therefore,
@@
Let $\ell_{1},\dots,\ell_{c}$ be the primes that divide $|G|$ and denote by $G(\ell_{i})$ the $\ell_{i}$-Sylow subgroup of $G$. Since $G$ is a nilpotent group, we have the product decomposition $G:=\prod_{i=1}^{c}G(\ell_{i})$. Denote by $\ell_{G}$ the smallest prime factor of $|G|$. Given $\sigma\in S_{n}$ and $g\in G$ we compute the index of $(\sigma,g)\in S_{n|G|}$ using the Lemma 2.4. The method employed by Wang for abelian groups, as detailed in [Wan21, Lemmas 2.5, 2.6 and 2.7] generalizes to nilpotent groups. For the sake of completeness, we provide details. In what follows, we set $G^{\ast}$ to denote $G\backslash\{1_{G}\}$ and $\operatorname{ind}(G):=\operatorname{min}\{\operatorname{ind}(g)\mid g\in G^{\ast}\}$.
@@
Let $G$ be a finite nilpotent group in its regular representation $G\hookrightarrow S_{|G|}$.
@@
    (2.1) $\mathrm{ind}((12),g)/|G|>2\text{ and }\mathrm{ind}((123),g)/|G|>1$
@@
    (2.2) $\mathrm{ind}((12),g)/|G|>2\text{ and }\mathrm{ind}((12)(34),g)/|G|>1,$
    (2.3) $\mathrm{ind}((123),g)/|G|>3\text{ and }\mathrm{ind}((1234),g)/|G|>2$
@@
    (2.4) $\mathrm{ind}(\sigma,g)/|G|\geq 1+\mathrm{ind}(\sigma)-1/7$
@@
    (2.5) $\mathrm{ind}(\sigma,g)/|G|\geq 12/7+\mathrm{ind}(\sigma).$
@@
We begin with (1). For $(12)\in S_{3}$, using Lemma 2.4 we deduce that for all $g\in G$ we have that $\mathrm{ind}((12),g)=|G|+2\cdot\mathrm{ind}(g)$ as $\mathrm{ind}(12)=1$. We deduce that
$\mathrm{ind}((12),g)/|G|\geq 1+2\cdot\mathrm{ind}(G)/|G|>2$
since $\ell_{G}>2$.
For $(123)\in S_{3}$, if $3\nmid|G|$, then again using Lemma 2.4 we obtain that for all $g\in G$, $\mathrm{ind}((123),g)=2|G|+\mathrm{ind}(g)$ as $\mathrm{ind}(123)=2$. Hence the inequality holds in this case. If $3\mid|G|$, then $G=G(3)\times\widetilde{G}$. Given any $g\in G$ we can write it as $(g(3),\widetilde{g})\in S_{|G(3)|}\times S_{\widetilde{G}}\subset S_{|G|}$. Further we note that $\mathrm{ind}((123),g)=\mathrm{ind}((123),(g(3),\widetilde{g}))=\mathrm{ind}(((123),g(3)),\widetilde{g})$ in $S_{|G|}$. Since $3\nmid|\widetilde{G}|$ we can use Lemma 2.4 to deduce that
$\mathrm{ind}(((123),g(3)),\widetilde{g})=\mathrm{ind}((123,g(3))\cdot(|\widetilde{G}|-\mathrm{ind}(\widetilde{g}))+3\cdot|G(3)|\cdot\mathrm{ind}(\widetilde{g}).$
Now, since the smallest possible value of $\mathrm{ind}((123,g(3))$ is $2|G(3)|$ when $g(3)$ is identity and smallest possible value for $\mathrm{ind}(\widetilde{g})$ is $\mathrm{ind}(\widetilde{G})$ by definition. Hence,
$\mathrm{ind}((123,g(3))\cdot(|\widetilde{G}|-\mathrm{ind}(\widetilde{g}))+3\cdot|G(3)|\cdot\mathrm{ind}(\widetilde{g})\geq 2|G|+|G(3)|\cdot\mathrm{ind}(\widetilde{G})>|G|.$
This concludes the proof.
@@
(3) Using Lemma 2.4 we conclude that
$\begin{split}\mathrm{ind}(\sigma,g)=&\mathrm{ind}(\sigma)|G|+5\cdot\mathrm{ind}(g)-\mathrm{ind}(\sigma)\cdot\mathrm{ind}(g)\\ \geq&\mathrm{ind}(\sigma)\cdot|G|+(5-\mathrm{ind}(\sigma))\cdot\mathrm{ind}(g)\geq\mathrm{ind}(\sigma)\cdot|G|+(5-\mathrm{ind}(\sigma))\cdot\mathrm{ind}(G)\\ \geq&\mathrm{ind}(\sigma)\cdot|G|+(5-\mathrm{ind}(\sigma))\cdot(\ell_{G}-1)|G|/\ell_{G}\\ \geq&\mathrm{ind}(\sigma)\cdot|G|+(5-\mathrm{ind}(\sigma))\cdot 6|G|/7\\ \geq&\operatorname{ind}(\sigma)|G|+6|G|/7,\end{split}$
since $\ell_{G}\geq 7$. Dividing both sides of the above inequality by $|G|$, we find that
$\frac{\operatorname{ind}(\sigma,g)}{|G|}\geq 1+\operatorname{ind}(\sigma)-\frac{1}{7},$
and this proves (2.4). For $\sigma\in S_{5}$ not conjugate to $(12345)$ we have that $(5-\mathrm{ind}(\sigma))\geq 2$ and hence
$\mathrm{ind}(\sigma)\cdot|G|+(5-\mathrm{ind}(\sigma))\cdot 6|G|/7\geq\mathrm{ind}(\sigma)\cdot|G|+12|G|/7,$
from which (2.5) follows. This concludes the proof of Proposition. ∎